OpenStudy (a1234):

Evaluate the two trigonometric functions using its period as an aid. 1. cos (9pi/2) = ______ Answer - 0 2. cos (-9pi/2) = ______ Answer - 0

OpenStudy (anonymous):

\[\cos \frac{ 9 \pi }{ 2 }=\cos \left( 4 \pi+\frac{ \pi }{ 2 } \right)=\cos \frac{ \pi }{ 2 }=0\] \[\cos \left( -\theta \right)=\cos\theta\ \]

OpenStudy (a1234):

So are both my answers correct?

OpenStudy (anonymous):

yes